# Electron-mediated anharmonicity and its role in the Raman spectrum of graphene

### v1

@@ Abstract
The Raman active G mode in graphene exhibits strong coupling to electrons, yet the comprehensive treatment of this interaction in the calculation of its temperature-dependent Raman spectrum remains incomplete. In this study, we calculate the temperature dependence of the G mode frequency and linewidth, and successfully explain the experimental trend, by accounting for the contributions arising from the first-order electron-phonon coupling, electron-mediated phonon-phonon coupling, and standard lattice anharmonicity. The generality of our approach enables its broad applicability to study phonon dynamics in materials where both electron-phonon coupling and anharmonicity are important.
@@ Introduction
Understanding phonons and their interactions is crucial for elucidating the role of lattice dynamics in material properties Giustino (2017), such as carrier and lattice transport, optical response, superconductivity, and structural phase transitions. Raman spectroscopy, a powerful and widely applicable experimental technique, plays a key role in studying temperature-dependent phonon behavior. This technique is particularly powerful for atomically thin materials like graphene and transition metal dichalcogenides Ferrari and Basko (2013); Zhang et al. (2015), where the position of the phonon peaks can reveal the actual doping concentration as well as the number of layers Das et al. (2008); Tan et al. (2011); Zhao et al. (2013). These experiments highlight the crucial role of electron-phonon coupling (EPC) and its competition with phonon-phonon interaction, showing how the interplay between phonons and electronic excitations influences the Raman features of layered materials Rafailov et al. (2002); Yoon et al. (2013); Sohier et al. (2019); Sarkar et al. (2020); Paul et al. (2021). In this regard, the case of graphene is especially intriguing Berciaud et al. (2010); Ferrante et al. (2018); Howard et al. (2011).
@@
One of the open questions in the Raman spectrum of graphene concerns the source of the observed temperature dependence for the G mode frequency and linewidth. From a theoretical point of view, four-phonon anharmonicity partially explains the observed G mode temperature trends in graphene and graphite Bonini et al. (2007); Han et al. (2022); Lin et al. (2011); Montagnac et al. (2013); Chen et al. (2021); Giura et al. (2012). On the other hand, Raman experiments done under ultrafast laser excitation, prove that electrons must be involved in the relaxation process in graphene Ferrante et al. (2018). This prompts us to question whether the G mode frequency decrease and an anomalous linewidth temperature behaviour could be the result of a combined effect of anharmonicity and EPC. In spite of its omnipresence, EPC is considered to have a weak contribution to the G mode linewidth at medium to strong doping: long-wavelength inter-band electron transitions are then prohibited by Pauli’s exclusion principle Ando (2006); Chae et al. (2010). Contrary to experimental observations Howard et al. (2011); Chen et al. (2023), this would imply that highly-doped graphene should have a zero EPC linewidth contribution, in spite of a strong doping-induced increase in overall EPC. For a long time it has been considered that the spectral signature of EPC should be a temperature-induced linewidth decrease Ponosov et al. (1998, 2005); Lazzeri and Mauri (2006); Bonini et al. (2007). However, this belief was often based on first-order calculations of EPC, which may not be adequate to capture the relevant physics in this case. A similar debate arose for magnesium diboride (MgB2).
@@
Similarities between MgB2 and graphene begin in their structure. Carbon atoms, arranged in a hexagonal pattern in graphene are effectively substituted with boron (B) atoms in MgB2. Meanwhile, magnesium atoms are positioned above and below the boron planes, introducing an intrinsic doping of the boron bands. The in-plane stretching of B-B bonds also gives rise to a phonon mode characterized by E2g symmetry. Due to the strong EPC, the E2g mode is responsible for a superconducting transition Bohnen et al. (2001); Kortus et al. (2001); Liu et al. (2001); Choi et al. (2002); Eiguren and Ambrosch-Draxl (2008); Calandra et al. (2010); Margine and Giustino (2013). Furthermore, due to its significant anharmonicity, the strongly temperature-dependent Raman spectra can only be explained outside the scope of the standard anharmonic theory, by employing higher-order electron-phonon scattering within the framework of the NA theory Cappelluti (2006); Saitta et al. (2008); Yildirim et al. (2001); Ponosov and Streltsov (2017). The corresponding first-principles theoretical advances beyond the standard NA theory have already been made for EPC in certain cases Novko (2018); Novko et al. (2016, 2018); Novko (2020). These amount to employing memory function methods originally derived to calculate intra-band relaxation processes in the dynamical conductivity Marsiglio and Carbotte (2008); Kupčić and Barišić (2005); Kupčić (2017); Novko (2017); Kupčić (2015); Allen (2015); Sharapov and Carbotte (2005); Chakraborty and Allen (1978), Raman response functions Kupčić (2008) or even exciton-phonon coupling Antonius and Louie (2022). In this way, one can account for the dynamical higher-order electron-phonon scattering, through the inclusion of a phonon-induced electron-hole pair self-energy by solving the Holstein problem Holstein (1964).
@@
The first-order intraband term is static and contains scattering processes within a narrow energy window around the chemical potential, broadened by temperature [see Eq. (2)]. It has no imaginary part, so it does not contribute to the EPC-induced linewidth. In graphene and other materials, even though the dynamical first-order intraband phonon self-energy contributions vanish in the long-wavelength limit, Raman experiments still suggest that EPC in fact must somehow contribute to the linewidth. In first-order perturbation theory, the excited phonon interacts with electrons via the generation of “bare” electron-hole pairs, while discarding the processes where the generated electron-hole pairs are renormalized by the rest of the system’s phonons.
@@
The factors in Eq. (Results and discussion), include the Eliashberg spectral function, the renormalized or quasi-particle DOS as well as the Fermi-Dirac f() and Bose-Einstein n) occupation factors. The same expression is commonly used for description of higher electron-phonon scattering terms in optical conductivity formula. The real part of the electron-hole self-energy can be obtained with the Kramers-Kronig transformation of .
@@
where energies are measured from the T=0 Fermi level. Equations (Results and discussion)-(2) need to be iterated self-consistently for every temperature until the temperature-dependent chemical potential and the corresponding quasi-particle DOS conserve the number of particles. Due to the complexity of the procedure, when calculating the chemical potential shifts, we resort to the on-shell approximation in the electron self-energy Eq. (Results and discussion). Note that the new value for the chemical potential enters the indirectly through the spectral function and the N0.
@@
In Fig. 3(a) we show the temperature variations in the imaginary part of the electron self-energy. At low temperatures, a characteristic energy dependence can be recognized resulting from the linear bands and the interaction with the optical phonons (i.e., G and A modes) Carbotte et al. (2010). The resulting renormalization of the quasi-particle DOS with temperature is shown in Fig. 3(b). The remaining panels in Fig. 3 show how the three different approaches to the calculation affect the electron-hole and phonon properties. The simplest approximation amounts to neglecting the energy dependence in the DOS. For graphene, it leads to almost no renormalization of the frequency, and weak linewidth contribution (blue line). On the other hand, considering the variations of the DOS with energy leads to larger values for the electron-hole self-energy, strong G mode renormalizations and significant linewidth contribution (red and yellow lines). The characteristic rapidly increasing DOS in graphene with meV, leads to the higher contribution coming from the ratio in Eq. (Results and discussion) (red), or for the case when quasi-particle DOS is included (yellow), which is then almost always larger than 1. Note how the proper inclusion of dynamical EPC leads to a quadratic temperature dependence of in the low temperature limit Allen and Silberglitt (1974); Girotto and Novko (2023a). Here we emphasize once again that to obtain the proper EPC-induced temperature-dependent phonon dynamics, it is crucial to account for the full self-consistent interactions of electron-hole pairs (inclusion of electron-hole self-energies with renormalized quasi-particle electronic structure) and conservation of the number of carriers (inclusion of interacting chemical potential shifts), both of which are usually omitted in the simulation of the phonon dynamics of Raman-active modes in graphene.
@@
If anharmonicity cannot be addressed perturbatively, it becomes challenging to relinquish the benefits of treating phonons as well-defined quasiparticles. Several methods have already been developed to treat this complexity, based on molecular dynamics (MD) Car and Parrinello (1985) (or path-integral MD if quantum effects in ionic motion are included Ceperley (1995)), or on a stochastic sampling of the potential energy surface. We compare these two approaches in our analysis of graphene’s Raman frequencies and linewidths.
The Temperature Dependent Effective Potential (TDEP) method extracts interatomic force constants (IFCs) by fitting displacements and forces directly from molecular dynamics (MD) simulations Knoop et al. (2024); Castellano et al. (2023); Hellman et al. (2011, 2013); Hellman and Abrikosov (2013) driven by forces from ab initio or Machine Learning (ML) calculations. This method intrinsically incorporates anharmonic effects both at the harmonic level, giving rise to an effective harmonic theory, and at higher orders (presently 3rd and 4th order IFCs), all of which are now explicitly temperature dependent. Direct sampling from MD allows for the system’s canonical ensemble to be unrestricted in its functional form, meaning it can be applied even at very high temperatures where anharmonicity is large. Quantum effects can be added by performing path integral MD simulations to sample the temperature dependent dynamics. Some of us have recently shown that the TDEP method results in an exact solution of the static part of the lattice dynamics, and therefore the exact static phonon frequencies Castellano et al. (2023).
The stochastic self-consistent harmonic approximation (SSCHA Monacelli et al. (2021); Errea et al. (2014); Bianco et al. (2017)), on the other hand, is based on the Gibbs–Bogoliubov free energy variational principle. Within SSCHA, the Gibbs-Bogoliubov free energy functional is minimized by employing trial harmonic potentials, and a stochastic approach is used to sample the potential energy surface according to the trial harmonic statistics. Once the minimization is performed and the free energy is obtained, the second-order IFCs, defined as the free energy Hessian with respect to the nuclei’s average positions, are also accessible. In SSCHA, both anharmonicity and quantum effects are included (the nuclear dynamics are treated at the quantum level, as the nuclear kinetic operator is fully accounted for), but the stochastic ensembles are constrained to be Gaussian.
@@
By using the SSCHA code, we performed the calculation of the SSCHA-phonon self-energy within the dynamical bubble approximation Monacelli et al. (2021), as shown diagramatically in Fig. 4(b). The triangular vertices indicate third-order SSCHA force constants, defined as the stochastic averages of the third-order derivatives of the potential with respect to the atomic positions. Within the TDEP framework, we computed similar phonon self-energy, but with both third and fourth order force constants.
The phonon spectral function is determined as the imaginary part of the phonon propagator. In this work, we compute it using the no-mode mixing approximation and further approximate it with a Lorentzian profile for each mode. We carry out calculations across a range of temperatures, then obtain the T=0 K frequency either by linearly extrapolating the temperature-dependent results or via molecular dynamics, excluding the zero-point frequency shift. This T=0 K frequency serves as a reference, enabling us to calculate the anharmonic phonon self-energy based on the frequency shift observed with temperature.
@@
The G mode frequency is found as the position of the peak of the calculated phonon spectral function Eq. (10). Depending on which self-energy contributions are included, we obtain the interband (i.e., first-order EPC), intraband (i.e., higher-order EPC), anharmonic, and total phonon renormalization with temperature. Note that our final results also incorporate the effects of the thermal expansion of the unit cell.
@@
Raman spectrum of graphene. In Fig. 5 we show the role of EPC in G mode Raman features for three different dopings in successive columns. If the Fermi level is 100 meV above the Dirac point, a G phonon with energy of meV directly facilitates the interband contribution to the EPC-induced phonon linewidth. In the second column, doping sets meV, where Pauli blocking should lead to negligible interband contributions. The third case corresponds to graphene doped up to the Van Hove singularity (VHS) in the DOS, which should approximately correspond to highly-doped intercalated graphene Howard et al. (2011); Chen et al. (2023). For the temperature dependence of both G mode frequency and linewidth we explicitly distinguish between the interband and intraband contributions to the phonon self-energy. Focusing first on the G mode frequency shift [see Figs. 5(a), 5(c), 5(e)] we notice that the intraband contribution grows with increased doping or higher temperatures. The interband contribution increases the G mode frequency with temperature until, for larger doping, its effect becomes negligible. Anharmonic effects are known to reduce the G mode frequency with an increase in temperature, especially the four-phonon contribution Lin et al. (2011); Bonini et al. (2007); Chen et al. (2021). Since the intraband contribution corresponds effectively to an electron-mediated phonon-phonon interaction Varma and Simons (1983); Novko et al. (2018), it shows a characteristic anharmonic-like increase with temperature. More striking contributions from the intraband phonon self-energy are observed for the G mode linewidth. Contrary to the common belief that the total EPC-induced linewidth derives from the interband phonon self-energy, we obtain a much larger intraband contribution, which has an opposite temperature trend. The weak initial linewidth decrease at low T, observed e.g. in Ref. 50; 44; 55 comes from the EPC interband term and is therefore present only when the chemical potential is smaller than the G mode frequency. The intraband contribution, on the other hand, increases with both T and doping. The total EPC-induced linewidth eventually turns out to be a solely intraband effect for moderately and highly-doped graphene.
@@
Overlooking higher-order electron-phonon scattering effects leads to an incomplete description of the Raman G mode spectral features. The EPC-induced changes of frequency and linewidth with temperature are almost entirely attributable to the commonly disregarded intraband contribution, and the experimentally observed behavior results from a cooperative interplay with anharmonicity. This supports the conclusion drawn in Ref. 12 that anharmonicity alone is not substantial enough to account for the full linewidth. Interestingly, our results for the highest doping case resembles the Raman spectroscopy experimental result from Ref. 51, where photo-doping is induced with lasers of high power. According to Berciaud 12 this is equivalent to large temperatures with . They report an unconventional decrease, followed by an increase in frequency, and an increase followed by a decrease in linewidth as a function of laser power (i.e., increase of temperature). The (rigid band) doping level used in our work, differs from the experimental one from Ref. 51, but the frequency and linewidth temperature dependence show very similar non-monotonous trends. The only mechanism able to explain such behaviour is higher-order EPC: the standard lattice-driven anharmonic contribution is constantly increasing. The same unconventional trend in the temperature dependence of the phonon linewidth was observed for mode in MgB2, and was explained with the EPC-induced anharmonicity Ponosov and Streltsov (2017); Novko et al. (2020).
@@
In the remaining two panels of Fig. 5, we show the doping dependence of the linewidth and inverse lifetime of electron-hole pairs at 300 K. In panel (g) we show how the intraband contribution to the linewidth becomes increasingly more important with doping in relative terms. This result also explains the decrease and subsequent increase of phonon linewidth , as a function of graphene doping with FeCl3Chen et al. (2023). The increase as a function of doping was explained in Ref. 57 in terms of electronic state broadening due to electron-electron scatterings in Raman process, however, here we show that the EPC-related scatterings could also explain this behavior. In panel Fig. 5(h) we show the linear dependence of the electron-hole inverse lifetime on doping. The slope of the curve amounts to 0.027, similar to the estimated value of 0.021 reported in Ref. 57 for intraband EPC-related transitions near .
@@
Having established the important role of the dynamical intraband contribution to the phonon self-energy at higher temperatures and dopings, we simulate the full Raman spectrum of graphene, with both the EPC and standard anharmonic effects obtained with SSCHA and TDEP. Note that here in the main text we report the TDEP anharmonic results obtained with stochastic sampling of the potential energy surface, i.e., sTDEP, while in the Supplementary Figure 2 we show and compare the TDEP results based on MD.
@@
Results are presented in Fig. 6 at a smaller temperature range than in Fig. 5. The chosen dopings are compared to Raman experiments done on pristine graphene and graphite in panels (a)-(b) Montagnac et al. (2013); Liu et al. (2019); Chae et al. (2010) and doped graphene on SiO2 in panels (c)-(d) Berciaud et al. (2010). Frequency dependence is largely determined by the anharmonic effects for both doping regimes, since the intraband EPC effects start contributing at larger temperatures. With both sTDEP and SSCHA we obtain a fairly good agreement with the temperature frequency trend observed in experiments for both cases.
@@
The importance of the proper inclusion of EPC effects along with the standard anharmonicity is more pronounced for the G mode linewidth. With SSCHA, only three-phonon contributions are included for the linewidths. In agreement with Ref. 49, we find that these three-phonon contributions are rather small, and even together with the much larger EPC effects the experimental linewidth values are still underestimated. In Supplementary Figures 3 and 4 we compare the three-phonon contributions as obtained with SSCHA and TDEP, and show that the best agreement is obtained when TDEP is combined with stochastic sampling of the potential energy surface. With TDEP we are able to account also for the four-phonon contribution that is an order of magnitude larger than the three-phonon term (see Supplementary Figures 3 and 4), leading to an excellent agreement with the experimental linewidths. In Supplementary Figures 5 and 6 we show that the thermal expansion significantly impacts the four-phonon contribution and is important to reach the final agreement with the experiments.

### v2

@@ Abstract
The Raman active G mode in graphene exhibits a strong coupling to electrons, yet the comprehensive treatment of this interaction in the calculation of its temperature-dependent Raman spectrum remains incomplete. In this study, we calculate the temperature dependence of the G-mode frequency and linewidth, and successfully explain the experimental trend, by accounting for the contributions arising from the first-order electron-phonon coupling, electron-mediated phonon-phonon coupling, and standard lattice anharmonicity. The generality of our approach enables its broad applicability to study phonon dynamics in materials where both electron-phonon coupling and anharmonicity are important.
@@ Introduction
Understanding phonons and their interactions is crucial to elucidate the role of lattice dynamics in material properties Giustino (2017), such as carrier and lattice transport, optical response, superconductivity, and structural phase transitions. Raman spectroscopy, a powerful and widely applicable experimental technique, plays a key role in studying temperature-dependent phonon behavior. This technique is particularly powerful for atomically thin materials like graphene and transition metal dichalcogenides Ferrari and Basko (2013); Zhang et al. (2015), where the position of the phonon peaks can reveal the actual doping concentration as well as the number of layers Das et al. (2008); Tan et al. (2011); Zhao et al. (2013). These experiments highlight the crucial role of electron-phonon coupling (EPC) and its competition with phonon-phonon interaction, showing how the interplay between phonons and electronic excitations influences the Raman features of layered materials Rafailov et al. (2002); Yoon et al. (2013); Sohier et al. (2019); Sarkar et al. (2020); Paul et al. (2021). In this regard, the case of graphene is especially intriguing Berciaud et al. (2010); Ferrante et al. (2018); Howard et al. (2011).
@@
One of the open questions in the Raman spectrum of graphene concerns the source of the observed temperature dependence for the G mode frequency and linewidth. From a theoretical point of view, four-phonon anharmonicity partially explains the observed G mode temperature trends in graphene and graphite Bonini et al. (2007); Han et al. (2022); Lin et al. (2011); Montagnac et al. (2013); Chen et al. (2021); Giura et al. (2012). On the other hand, Raman experiments done under ultrafast laser excitation, prove that electrons must be involved in the relaxation process in graphene Ferrante et al. (2018). This prompts us to question whether the G mode frequency decrease and an anomalous linewidth temperature behaviour could be the result of a combined effect of anharmonicity and EPC. In spite of its omnipresence, EPC is considered to have a weak contribution to the G mode linewidth at medium to strong doping: long-wavelength interband electron transitions are then prohibited by Pauli’s exclusion principle Ando (2006); Chae et al. (2010). Contrary to experimental observations Howard et al. (2011); Chen et al. (2023), this would imply that highly-doped graphene should have a zero EPC linewidth contribution, in spite of a strong doping-induced increase in overall EPC. For a long time it has been considered that the spectral signature of EPC should be a temperature-induced linewidth decrease Ponosov et al. (1998, 2005); Lazzeri and Mauri (2006); Bonini et al. (2007). However, this belief was often based on first-order calculations of EPC, which may not be adequate to capture the relevant physics in this case. A similar debate arose for MgB2.
@@
The similarities between MgB2 and graphene begin in their structure. Carbon atoms, arranged in a hexagonal pattern in graphene are effectively substituted with boron atoms in MgB2. Meanwhile, magnesium atoms are positioned above and below the boron planes, introducing an intrinsic doping of the boron bands. The in-plane stretching of B-B bonds also gives rise to a phonon mode characterized by E2g symmetry. Due to the strong EPC, the E2g mode is responsible for a superconducting transition Bohnen et al. (2001); Kortus et al. (2001); Liu et al. (2001); Choi et al. (2002); Eiguren and Ambrosch-Draxl (2008); Calandra et al. (2010); Margine and Giustino (2013). Furthermore, due to its significant anharmonicity, the strongly temperature-dependent Raman spectra can only be explained outside the scope of the standard anharmonic theory, by employing higher-order electron-phonon scattering within the framework of the NA theory Cappelluti (2006); Saitta et al. (2008); Yildirim et al. (2001); Ponosov and Streltsov (2017). The corresponding first-principles theoretical advances beyond the standard NA theory have already been made for EPC in certain cases Novko (2018); Novko et al. (2016, 2018); Novko (2020). These amount to employing memory function methods originally derived to calculate intraband relaxation processes in the dynamical conductivity Marsiglio and Carbotte (2008); Kupčić and Barišić (2005); Kupčić (2017); Novko (2017); Kupčić (2015); Allen (2015); Sharapov and Carbotte (2005); Chakraborty and Allen (1978), Raman response functions Kupčić (2008) or even exciton-phonon coupling Antonius and Louie (2022). In this way, one can account for the dynamical higher-order electron-phonon scattering, through the inclusion of a phonon-induced electron-hole pair self-energy by solving the Holstein problem Holstein (1964).
@@ Results
The first-order intraband term is static and contains scattering processes within a narrow energy window around the chemical potential, broadened by temperature [see Eq. (2)]. It has no imaginary part, so it does not contribute to the EPC-induced linewidth. In graphene and other materials, even though the dynamical first-order intraband phonon self-energy contributions vanish in the long-wavelength limit, Raman experiments still suggest that EPC in fact must somehow contribute to the linewidth. In the first-order perturbation theory, the excited phonon interacts with electrons via the generation of “bare” electron-hole pairs, while the processes where the generated electron-hole pairs are renormalized by the rest of the system’s phonons are discarded.
@@
The factors in Eq. (2), include the Eliashberg spectral function, the renormalized or quasi-particle DOS as well as the Fermi-Dirac f() and Bose-Einstein n) occupation factors. The same expression was used for description of higher electron-phonon scattering terms in optical conductivity formula Sharapov and Carbotte (2005). The real part of the electron-hole self-energy can be obtained with the Kramers-Kronig transformation of .
@@
where energies are measured from the T=0 Fermi level. Equations (2)-(3) need to be iterated self-consistently for every temperature until the temperature-dependent chemical potential and the corresponding quasi-particle DOS conserve the number of particles. Due to the complexity of the procedure, when calculating the chemical potential shifts, we resort to the on-shell approximation in the electron self-energy Eq. (2) where the general energy argument is substituted by the Kohn-Sham value Giustino (2017). Note that the new value for the chemical potential enters the indirectly through the spectral function and the N0.
@@
In Fig. 3(a) we show the temperature variations in the imaginary part of the electron self-energy. At low temperatures, a characteristic energy dependence can be recognized resulting from the linear bands and the interaction with the optical phonons (i.e., G and A modes) Carbotte et al. (2010); Park et al. (2007). The resulting renormalization of the quasi-particle DOS with temperature is shown in Fig. 3(b). The remaining panels in Fig. 3 show how the three different approaches to the calculation affect the electron-hole and phonon properties. The simplest approximation amounts to neglecting the energy dependence in the DOS. For graphene, it leads to almost no renormalization of the frequency, and weak linewidth contribution (blue line). On the other hand, considering the variations of the DOS with energy leads to larger values for the electron-hole self-energy, strong G mode renormalizations and significant linewidth contribution (red and yellow lines). The characteristic rapidly increasing DOS in graphene with meV, leads to the higher contribution coming from the ratio in Eq. (2) (red), or for the case when quasi-particle DOS is included (yellow), which is then almost always larger than 1. Note how the proper inclusion of dynamical EPC leads to a quadratic temperature dependence of in the low temperature limit Allen and Silberglitt (1974); Girotto and Novko (2023a). Here we emphasize once again that to obtain the proper EPC-induced temperature-dependent phonon dynamics, it is crucial to account for the full self-consistent interactions of electron-hole pairs (inclusion of electron-hole self-energies with renormalized quasi-particle electronic structure) and conservation of the number of carriers (inclusion of interacting chemical potential shifts), both of which are usually omitted in the simulation of the phonon dynamics of Raman-active modes in graphene, as well as in other systems where EPC plays an important role. Using the model Hamiltonians, this electron-mediated anharmonicity was argued to be important for charge-density-wave (CDW) materials where EPC is strong and these higher-order effects contribute to the CDW transition temperature and renormalization of electronic structure Varma and Simons (1983); Flicker and van Wezel (2015); Yoshiyama et al. (1986).
@@
Further, it is important to stress that we use a standard semi-local DFT functional in order to calculate the EPC properties for graphene, while it was shown that beyond-DFT corrections to electron correlations (e.g., GW approximation and hybrid functionals) modify the electronic structure, increase the Fermi velocity around the Dirac points Trevisanutto et al. (2008), and enhance the EPC strengths of both G and A phonon modes Lazzeri et al. (2008); Attaccalite et al. (2010). These beyond-DFT corrections might impact the evaluation of the intraband phonon self-energy, however, it is not clear to what extent, considering that the increase of the Fermi velocity will decrease the value of and in Eq. (8), while the increase of the EPC strength will increase and . The competition of these electron correlation effects in the evaluation of the phonon self-energy is an interesting problem that we also leave for the future work.
@@
If anharmonicity cannot be addressed perturbatively, it becomes challenging to relinquish the benefits of treating phonons as well-defined quasiparticles. Several methods have already been developed to treat this complexity, based on molecular dynamics (MD) Car and Parrinello (1985) (or path-integral MD if quantum effects in ionic motion are included Ceperley (1995)), or on a stochastic sampling of the potential energy surface. We mostly focus on the latter in our analysis of graphene’s Raman frequencies and linewidths.
The Temperature Dependent Effective Potential (TDEP) method extracts interatomic force constants (IFCs) by fitting displacements and forces directly from molecular dynamics (MD) simulations Knoop et al. (2024); Castellano et al. (2023); Hellman et al. (2011, 2013); Hellman and Abrikosov (2013) driven by forces from ab initio or Machine Learning (ML) calculations. This method intrinsically incorporates anharmonic effects both at the harmonic level, giving rise to an effective harmonic theory, and at higher orders (presently 3rd and 4th order IFCs), all of which are now explicitly temperature dependent.
The stochastic self-consistent harmonic approximation (SSCHA Monacelli et al. (2021); Errea et al. (2014); Bianco et al. (2017)), on the other hand, is based on the Gibbs–Bogoliubov free energy variational principle. Within SSCHA, the Gibbs-Bogoliubov free energy functional is minimized by employing trial harmonic potentials, and a stochastic approach is used to sample the potential energy surface according to the trial harmonic statistics. SSCHA inherently accounts for both anharmonicity—incorporating contributions from all even orders of the potential energy surface—and quantum effects, as nuclear dynamics are fully treated at the quantum level, with the nuclear kinetic operator explicitly taken into account. However, the stochastic ensembles are constrained to be Gaussian. Once the minimization is performed and the free energy is obtained, an effective set of second-order interatomic force constants (IFCs) naturally emerges, defining the stochastic self-consistent harmonic (SCHA) noninteracting quasiparticles. The TDEP code has implemented an equivalent method using stochastic sampling, but with its own IFC fitting approach. This method is commonly known as sTDEP Shulumba et al. (2017) and in this work we compare it with the SSCHA method.
@@
Using the SSCHA code, we calculated the SCHA-phonon self-energy, explicitly accounting for interactions between SCHA phonons. We restricted our analysis to the so-called bubble approximation, shown diagrammatically in Fig. 4(b), where only third-order SCHA interaction vertices are included. These vertices are defined as the stochastic averages of the third-order derivatives of the potential energy surface with respect to atomic positions. In our evaluation of the SCHA-phonon self-energy, we did not include fourth-order SCHA interaction vertices Monacelli et al. (2021). Within the TDEP framework, we computed a similar phonon self-energy but including both third- and fourth-order interaction vertices between the auxiliary quasiparticles.
The phonon spectral function is determined as the imaginary part of the phonon propagator. In this work, we compute it using the no-mode mixing approximation and further approximate it with a Lorentzian profile for each mode. We carry out calculations across a range of temperatures, then obtain the K frequency either by linearly extrapolating the temperature-dependent results or via molecular dynamics, excluding the zero-point frequency shift. This K frequency serves as a reference, enabling us to calculate the anharmonic phonon self-energy based on the frequency shift observed with temperature.
@@
The G mode frequency is found as the position of the peak of the calculated phonon spectral function Eq. (10). Depending on which self-energy contributions are included, we obtain the interband (i.e., first-order EPC), intraband (i.e., higher-order EPC), anharmonic, and total phonon renormalization with temperature. Note that our final SSCHA and sTDEP results also incorporate the effects of the thermal expansion of the unit cell calculated using a quantum mechanical approach. The DFPT calculations and EPC properties do not include the effect of volume expansion.
@@
Raman spectrum of graphene. In Fig. 5 we show the role of EPC in G mode Raman features for three different dopings in successive columns. If the Fermi level is 100 meV above the Dirac point, a G phonon with energy of meV directly facilitates the interband contribution to the EPC-induced phonon linewidth. In the second column, doping sets meV, where Pauli blocking should lead to negligible interband contributions. The third case corresponds to graphene doped up to the Van Hove singularity (VHS) in the DOS, which should approximately correspond to highly-doped intercalated graphene Howard et al. (2011); Chen et al. (2023). For the temperature dependence of both G mode frequency and linewidth we explicitly distinguish between the interband and intraband contributions to the phonon self-energy. Focusing first on the G mode frequency shift [see Figs. 5(a), 5(c), 5(e)] we notice that the intraband contribution grows with increased doping or higher temperatures. The interband contribution increases the G mode frequency with temperature until, for larger doping, its effect becomes negligible. Anharmonic effects are known to reduce the G mode frequency with an increase in temperature, especially the four-phonon contribution Lin et al. (2011); Bonini et al. (2007); Chen et al. (2021). Since the intraband contribution corresponds effectively to an electron-mediated phonon-phonon interaction Varma and Simons (1983); Novko et al. (2018), it shows a characteristic anharmonic-like increase with temperature. More striking contributions from the intraband phonon self-energy are observed for the G mode linewidth. Contrary to the common belief that the total EPC-induced linewidth comes from the interband phonon self-energy, we obtain a much larger intraband contribution, which has an opposite temperature trend. The weak initial linewidth decrease at low T, observed e.g. in Ref. 50; 44; 55 comes from the EPC interband term and is therefore present only when the chemical potential is smaller than the half of the G mode frequency. The intraband contribution, on the other hand, increases with both T and doping. The total EPC-induced linewidth eventually turns out to be a solely intraband effect for moderately and highly-doped graphene.
@@
Overlooking higher-order electron-phonon scattering effects leads to an incomplete description of the Raman G mode spectral features. The EPC-induced changes of frequency and linewidth with temperature are almost entirely attributable to the commonly disregarded intraband contribution, and the experimentally observed behavior results from a cooperative interplay with anharmonicity. This supports the conclusion drawn in Ref. 12 that anharmonicity alone is not substantial enough to account for the full linewidth. Interestingly, our results for the highest doping case resembles the Raman spectroscopy experimental result from Ref. 51, where photo-doping is induced with lasers of high power. According to Berciaud Berciaud et al. (2010) this is equivalent to large temperatures with . They report an unconventional decrease, followed by an increase in frequency, and an increase followed by a decrease in linewidth as a function of laser power (i.e., increase of temperature). The (rigid band) doping level used in our work, differs from the experimental one from Ref. 51, but the frequency and linewidth temperature dependence show very similar non-monotonous trends. The only mechanism able to explain such behaviour is higher-order EPC: the standard lattice-driven anharmonic contribution is constantly increasing. The same unconventional trend in the temperature dependence of the phonon linewidth was observed for mode in MgB2, and was explained with the EPC-induced anharmonicity Ponosov and Streltsov (2017); Novko et al. (2020).
@@
In the remaining two panels of Fig. 5, we show the doping dependence of the linewidth and inverse lifetime of electron-hole pairs at 300 K. In panel (g) we show how the intraband contribution to the linewidth becomes increasingly more important with doping in relative terms. This result also explains the decrease and subsequent increase of phonon linewidth , as a function of graphene doping with FeCl3Chen et al. (2023). The increase as a function of doping was explained in Ref. 57 in terms of electronic state broadening due to electron-electron scatterings in Raman process, however, here we show that the EPC-related scatterings could also explain this behavior. In Fig. 5(h) we show the linear dependence of the electron-hole inverse lifetime on doping. The slope of the curve amounts to 0.027, similar to the estimated value of 0.021 reported in Ref. 57 for intraband EPC-related transitions near .
@@
Having established the important role of the dynamical intraband contribution to the phonon self-energy at higher temperatures and dopings, we simulate the full Raman spectrum of graphene, with both the EPC and standard anharmonic effects obtained with SSCHA and TDEP. Note that here in the main text we report the TDEP anharmonic results obtained with stochastic sampling of the potential energy surface, i.e., sTDEP, while in the Supplementary Figure 2 we show and compare the TDEP results based on MD and PiMD.
@@
Results are presented in Fig. 6 at a smaller temperature range than in Fig. 5. The chosen dopings are compared to Raman experiments done on pristine graphene and graphite in panels (a)-(b) Montagnac et al. (2013); Liu et al. (2019); Chae et al. (2010) and doped graphene on SiO2 in panels (c)-(d) Berciaud et al. (2010). Frequency dependence is largely determined by the anharmonic effects for both doping regimes, since the intraband EPC effects start contributing at larger temperatures. With both sTDEP and SSCHA we obtain a fairly good agreement with the temperature frequency trend observed in experiments for both cases. The SSCHA and sTDEP results greatly overlap, proving the equivalence of the two methods. Note that the full results, which include both EPC effects and standard anharmonicity, slightly overestimate the experimental values for the undoped graphene (by meV), however, it is important to observe that the difference between our results and experiments is within the same scale as the difference between various experimental results. Moreover, the shift of our result with respoct to the experimental values can also be attributed to the fact that our DFPT calculation does not account for lattice thermal expansion. Lattice expansion would lead to a slight reduction of the phonon frequencies.
@@
The importance of the proper inclusion of EPC effects along with the standard anharmonicity is more pronounced for the G mode linewidth. With SSCHA, only three-phonon contributions are included for the linewidths. In agreement with Ref. 49, we find that these three-phonon contributions are rather small, and even together with the much larger EPC effects the experimental linewidth values are still underestimated. In Supplementary Figure 3 we show that the three-phonon contributions as obtained with SSCHA and sTDEP are quite similar. With TDEP we are able to account also for the four-phonon contribution that is an order of magnitude larger than the three-phonon term (see Supplementary Figures 3 and 4), leading to an excellent agreement with the experimental linewidths. This may raise the question of the importance of even higher orders. However, the larger magnitude of four-phonon compared to three-phonon terms is not due to a higher energy contribution of fourth-order anharmonicity but to the mirror-plane symmetry of graphene which forbids any interaction involving an odd number of flexural phonon. This sum rule restricts drastically the number of available three-phonon processes, leading to its lower contribution, while four-phonon term are less affected. Since forces predicted by the anharmonic potential up to fourth-order allows for an accurate reproduction of the full anharmonic forces, as shown in Supplementary Figure 5, we should expect higher order terms to be mainly negligible. In Supplementary Figures 4 and 6 we show that the inclusion of the quantum thermal expansion significantly impacts the four-phonon contribution and is important to reach the final agreement with the experiments. A classical approach in the volume expansion calculation leads to an overestimation of the phonon linewidth.